# Strategic Investment to Mitigate Transition Risks

### v1

@@ A Total Asset
$I$ Carbon Intensity
Table 1: Notations Description Table
@@ 2.3 Optimal Transition Investment Ratio without Government Intervention
The transition towards a low-carbon economy presents a complex challenge for enterprises, acting as a double-edged sword with respect to both arising opportunities and costs. While embracing low-carbon transformation holds the promise of environmental sustainability and enhanced competitiveness, it also entails a substantial amount of financial investments and operational adjustments. For many enterprises, particularly those operating in carbon-intensive industries, the costs associated with transitioning to cleaner technologies and sustainable practices can be daunting. These costs may include investments in renewable energy infrastructure, efficiency upgrades, and emission reduction measures, all of which require a substantial amount of capital expenditure and major operational restructuring in order to embark on the transition path.
As a result, some companies may hesitate to fully commit to low-carbon transformation, fearing the resulting financial burden and potential disruptions to their existing business models. This reluctant proclivity to adopt green practices could undermine a firm’s progress towards sustainability goals and exacerbate environmental challenges. In such cases, the role of government intervention becomes relevant in incentivizing and regulating enterprises’ low-carbon transformation efforts.
@@
Hence, this section aims at analyzing and delineating the optimal low-carbon transition investment strategies for companies, considering both scenarios with and without external government intervention. By comparing these two approaches separately, we hope to elucidate the efficacy and implications of different pathways toward sustainable transformation within the corporate landscape and to provide valuable insights and suggestions for policymakers, such as government entities, to better support and incentivize effective low-carbon strategies in the corporate sector.
2.3.1 Model Building
Maximizing profit holds a paramount position among the objectives of any company. This strategic action emanates from the foundational principle that profit arises from the difference between total revenue and total costs. Therefore, to effectively navigate and optimize profit margins, it becomes indispensable to dissect the constituents of both total revenue and total costs.
@@ 2.3.1 Model Building
In our framework, total revenue is represented by the unit selling price per unit of the product (denoted as $p$) multiplied by the total production volume. The total production volume is represented by $2\cdot\text{Sigmoid}(\alpha k)\cdot[(1-\alpha)\cdot A]^{\beta}$, where $\operatorname{Sigmoid}(\alpha k)=\frac{1}{1+e^{-\alpha k}}$. This formulation integrates several key parameters: the transition investment ratio ($\alpha$), indicating the allocation of total assets towards low-carbon initiatives; the enterprise low-carbon production efficiency coefficient ($k$), denoting the effectiveness of transitioning to cleaner technologies; and the original productivity coefficient ($\beta$), reflecting the baseline production efficiency. The sigmoid function encapsulates the joint impact of transition investments and the firm’s production efficiency on production volume, ensuring a realistic representation of the firm’s saturation effects. Meanwhile, the term $(1-\alpha)\cdot A$ captures the portion of the firm’s total assets available for production activities, adjusted by the original productivity coefficient $\beta$ to reflect the company’s inherent efficiency before investing in the low-carbon transition. Thus, the term $[(1-\alpha)\cdot A]^{\beta}$ represents the firm’s total production volume before investing in a low-carbon transition. Simultaneously, the firm allocates a portion $\alpha$ of its total assets $A$ towards the low-carbon transition investment. This investment is expected to enhance production efficiency, leading to an overall increase in production volume by a factor of $2\cdot\operatorname{Sigmoid}(\alpha k)$. By combining these various elements, the formula provides a greater understanding of how strategic decisions regarding transition investments and production efficiency influence the overall production capacity of the company.
@@
The design of this production function originates from a Cobb-Douglas production function, a widely used model in microeconomics that represents the relationship between inputs (typically capital and labor) and output in a production process. The general form of the Cobb-Douglas function is $Q=A\cdot K^{\beta}\cdot L^{1-\beta}$, where $Q$ is the total output, $A$ is a scaling or efficiency parameter, $K$ is capital input, $L$ is labor input, and $\beta$ is the elasticity of output with respect to capital. In our framework, the term $[(1-\alpha)\cdot A]^{\beta}$ corresponds to $K^{\beta}$, where $K$ represents capital. The sigmoid function, multiplied by 2, acts as an adaptive transformation of the efficiency parameter $A$ in Cobb-Douglas, reflecting the transition investment’s impact on production efficiency. Unlike the traditional Cobb-Douglas function, which often considers both capital and labor inputs, our model excludes the labor component due to the specific context of low-carbon investment decisions. This adaptation ensures that the production process incorporates diminishing marginal returns to transition investments ( $\alpha$ ) and captures the saturation effects as efficiency reaches its practical limits. Notably, we introduce a scaling factor of ‘2’ at the beginning of the sigmoid function. This adjustment is made primarily to ensure that the sigmoid’s baseline value remains at least 0.5 when the input ($\alpha k$) exceeds 0 in our model. By multiplying the sigmoid by ‘2’, we normalize its baseline rate to 1, resulting in a more pronounced increase above 1 as the values of $\alpha$ (Transition Investment Ratio) or $k$ (Enterprise Low-carbon Production Efficiency Coefficient) rise. This modification aligns with our prior expectations about this relationship.
@@
Regarding robustness, our results are likely to remain valid with alternative production functions such as Cobb-Douglas or CES (Constant Elasticity of Substitution), as these also exhibit concavity and diminishing marginal returns, which are key economic properties captured by our sigmoid-based framework. The trade-offs between transition investments and production efficiency, as well as the insights on profitability, depend on these shared characteristics and are not restricted to the functional form that we use for our study. However, CES functions, which allow for variable elasticity of substitution between inputs, could add an additional degree of flexibility and provide further insights into how substitutability between capital and efficiency affects production outcomes. While closed-form general solutions may be challenging with alternative production functions due to their increased complexity, numerical simulations could be in principle applied to extend our model to other functional forms. This flexibility demonstrates the potential for generalizing our framework while still maintaining the principle notions of diminishing returns and efficient allocation of low-carbon investments.
Based on the descriptions provided above, we can formulate the expression for ‘Revenue’ in the framework as follows:
[…]
@@
Next, we address the ‘Cost’ component of the framework, which comprises both the ‘Production Cost’ and the ‘Carbon Cost’. Denote ‘c’ as the production cost per unit. The production cost can then be expressed as ‘c’ multiplied by the unit amount (equivalent to the unit amount in the revenue part). Additionally, we incorporate the carbon cost in our framework, which is calculated as the product of the ‘Carbon Intensity’, the unit amount, and the ‘Carbon Price’. Here, the ‘Carbon Intensity’ is defined as $\left(2-\frac{2}{1+e^{-\alpha k}}\right)$, a value ranging from 0 to 1. When $\alpha k$ is zero (signifying no transition investment), the carbon intensity reaches its highest level of 1. Conversely, as transition investment increases, the carbon intensity decreases, indicating a reduction in emissions due to greater transition efforts. As mentioned, we define ‘Carbon Intensity’ as the ‘Carbon Emission’ divided by the total production volume, aligning with the traditional understanding of ‘Carbon Intensity’ discussed earlier. Based on these considerations, we can formulate the ‘Cost’ component in the framework as follows:
[…]
@@
To determine the optimal value of $\alpha$ under various conditions, we begin by taking the first-order derivative of the total profit with respect to $\alpha$. This yields a necessary first-order condition (FOC) as follows:
@@
This equation helps us identify potential critical points of maximizing the Equation (2.3). To further analyze these points, specifically to assess their convexity and determine whether they are minima, maxima, or saddle points arising from the formulation of the model, we compute the second-order derivative with respect to $\alpha$. The resulting second-order condition (SOC) is used to evaluate the nature of these critical points.
@@
From the analysis of the second derivative, we observe that the value is less than 0 when $\alpha$ and $\beta$ are both within the range of (0, 1). Consequently, the optimal value of $\alpha$ maximizes profit, aligning with the expectations set by the problem.
Our first objective is to understand how changes in specific parameters affect the optimal transition investment ratio. For instance, we aim at analyzing how the optimal value of $\alpha$ adjusts when the carbon price ($B$) increases from 0 to 2. This analysis will elucidate the impact of carbon pricing on the optimal transition investment ratio for the firm. However, explicitly expressing $\alpha$ in terms of other parameters analytically is challenging based on the FOC due to several inherent complexities. The equation features exponential terms such as $e^{\alpha(-k)}$, which complicate the isolation of $\alpha$ since these terms do not simplify easily into a linear or polynomial form. Additionally, the equation includes a combination of linear and non-linear terms involving multiple parameters ($A$, $B$, $c$, $p$, $\beta$, and $k$), further complicating the process. The presence of $\alpha$ in both the numerator and the denominator within the fractional expressions adds another layer of difficulty. Therefore, we will resort to numerical methods to demonstrate how the optimal transition investment ratio of the firm varies in response to incremental parameter changes.
2.3.2 Numerical Study
In this subsection, we conduct a detailed numerical analysis of each critical parameter affecting the optimal transition investment ratio ($\alpha$). Recall the optimal transition investment ratio represents the proportion of firm’s financial resources allocated towards its low-carbon initiatives relative to its total investment. A key variable is a specific variable within our framework that we systematically vary to observe its impact on the optimal $\alpha$. By focusing on one variable at a time, we can isolate its effect and gain deeper insights into its role within the overall system. This is known in Economics as a comparative static exercise.
While holding all other variables constant, we determine the optimal value of $\alpha$ for a range of values of a specific key variable. We then graphically represent the relationship between this key variable and the optimal $\alpha$. Additionally, we will present the maximum profit achievable for different values of the key variable. This approach allows us to visualize how changes in one variable influence both the optimal investment decision and the profitability of the company.
The baseline values of the variables in the framework are provided in Table 2. These values are set within a reasonable range for illustrative purposes and can be adjusted by readers interested in exploring scenarios across different industries. Building on this foundation, we will systematically vary one variable at a time to different values and analyze the resultant changes in the value of optimal $\alpha$.
@@ 2.3.2 Numerical Study
$p-c$ Selling Profit per unit 2
@@ A Total Asset 100
Table 2: Baseline Parameters for Analysis
From Figure 1 to Figure 4, the left plot illustrates how changes in the key variable affect the optimal transition investment ratio, aiming to maximize profit under each scenario. The right plot shows the corresponding maximum profit achieved under the optimal transition investment as the key variable changes. We will explain each figure in detail below.
Relationship Between Profit Margin ($p-c$) and Optimal Alpha (Figure 1)
Figure 1 reveals a distinctive pattern where the optimal value of alpha for the firm remains high and constant initially, even when the selling profit per unit is very low or negative, suggesting a strategic focus on substantial investments in machinery upgrades or efficiency improvements during periods of low profitability. This high investment level, indicated by a stable high value of alpha, which is likely to reflect a counterintuitive but proactive approach: companies invest heavily during financially challenging times to enhance production efficiency and reduce operational costs, aiming at setting the groundwork for future competitiveness and sustainability.
As the selling profit per unit begins to increase for the firm, the optimal value of alpha sharply declines, indicating that further heavy investments in transformation by the firm are no longer necessary. The decrease in the value of alpha suggests that the earlier investments by the firm have begun to yield improved operational efficiencies and higher profits, reducing the need for continued high expenditure in transformations. This strategic shift emphasizes a dynamic financial strategy where initial heavy investments by the firm during low-profit periods are scaled back as profitability improves, aligning the investment levels of the firm with the diminishing marginal benefits of further transformations.
Figure 1: Optimal Level of Alpha and Profit Graph w.r.t. $p-c$
Effect of Increasing Carbon Prices ($B$) on Optimal Alpha (Figure 2)
As shown in Figure 2, when the carbon pricing variable $B$ increases, it raises operational costs for the firm due to higher carbon taxes. In response, the optimal investment ratio $\alpha$ also increases, as shown in the left plot. This adjustment indicates a strategic move by the firm to allocate more resources towards low-carbon transition initiatives, which enhances production efficiency and reduces carbon emissions, thereby helping to offset the financial burden imposed by the carbon tax.
@@
Impact of Efficiency Parameter ($k$) on Optimal Alpha and Profit (Figure 3)
Combining insights from both plots in Figure 3, we can observe that as the efficiency parameter $k$ increases, profit consistently rises, while the optimal investment ratio $\alpha$ peaks around $k\approx 3.1$ and then gradually declines. This suggests that although higher $k$ values are beneficial for the company’s profitability, it does not require continuously increasing the investment ratio in low-carbon initiatives. Initially, when $k$ is low, the marginal benefit from efficiency gains is high, encouraging a higher allocation to low-carbon investments. However, as $k$ grows, the marginal efficiency gains diminish, making it less necessary to maintain a high $\alpha$. At higher $k$ values, profits continue to increase due to the enhanced production efficiency achieved by prior investments, allowing the company to reduce the investment ratio $\alpha$ while still benefiting from increased profits.
@@
Influence of Original Productivity Coefficient ($\beta$) on Optimal Alpha and Profit (Figure 4)
The plots in Figure 4 demonstrate that as the Original Productivity Coefficient ($\beta$) increases, profit grows exponentially while the optimal transition investment ratio ($\alpha$) declines. A higher $\beta$ indicates strong baseline productivity, reducing the need for extensive low-carbon investments since the firm’s production efficiency is already high. This makes the additional benefits from low-carbon investments less significant, allowing the firm to allocate fewer resources to $\alpha$. The exponential growth in profit with rising $\beta$ highlights the significant impact of inherent productivity on profitability, as firms with high baseline efficiency can achieve substantial profits without heavy investment in low-carbon initiatives. The decrease in optimal $\alpha$ reflects diminishing marginal returns on low-carbon investments for high-$\beta$ firms, where the benefits of further investment in $\alpha$ are minimal.
@@
2.3.3 Sensitivity Analysis
The sensitivity analysis was conducted to evaluate the impact of variations in critical parameters on the optimal alpha, maximum objective value, and profit change rate. The parameters scrutinized include the enterprise low-carbon production efficiency coefficient ($k$), production cost per unit ($c$), original productivity coefficient ($\beta$), and carbon price per unit ($B$). Adjustments for each parameter were performed at ±10% and ±50% based on the baseline values in Table 2, except for $\beta$, where the upper limit was constrained to 1.00 due to inherent limitations. The results are concluded in the Table 3 and the analysis of each parameter is presented below.
@@ 2.3.3 Sensitivity Analysis
Profit Change Rate (%) 60.249582 11.842530 $0$11.69191 -56.05747
@@
Alpha Change Rate (%) 54.473780 9.273146 $0$8.725114 -20.955217
@@
Profit Change Rate (%) 26.669326 4.834262 $0$4.600030 -20.871997
@@
The sensitivity analysis for $k$ reveals that as $k$ increases from 1.0 to 3.0, the optimal alpha rises from 0.14 to 0.41, which indicates that improving a firm’s low-carbon production efficiency encourages a larger investment by the firm in low-carbon technologies. This trend is further reflected in the maximum objective value, which grows from 40.39 to 62.52, showing that better firm’s efficiency leads to higher profitability for the firm. These results highlight a strong connection between firm’s low-carbon production efficiency and its profitability.
[…]
Production Cost Per Unit (c):
@@
As $c$ increases from 0.80 to 2.40, the optimal alpha rises from 0.30 to 0.52, suggesting that firms must invest more to counterbalance higher production costs. However, this comes at the cost of its profitability, as the maximum objective value drops sharply from 80.82 to 22.16, showing that rising production costs significantly undermine a company’s profitability potential.
@@
The alpha change rate starts at $-19.88\%$ for $c=0.80$, meaning that lower production costs lead to less investment in low-carbon technology. As $c$ rises, the alpha change rate becomes positive (38.13%), showing that companies need to increase their investments to cope with higher costs. Profit change rates, meanwhile, vary drastically, reaching a peak at $60.25\%$ for $c=0.80$, followed by a steep decline to $-56.06\%$ at $c=2.40$. This result illustrates the sensitivity of firm’s profitability to changes in its production costs and highlights that controlling production costs is crucial for firms whose goal is to sustain profits while investing in low-carbon solutions. If its production costs rise unchecked, the firm may struggle to maintain its profitability, even with optimal investments.
Original Productivity Coefficient $(\beta)$ :
@@
When it comes to the value of $\beta$, the sensitivity analysis overall shows a unique pattern: as $\beta$ increases from 0.40 to 1.00, the optimal alpha decreases from 0.58 to 0.30. This indicates that as its productivity improves, the firm requires less investment in low-carbon technologies to maintain profitability. Simultaneously, the maximum objective value surges from 10.34 to 116.73, highlighting the substantial impact of productivity on overall profitability of the firm. This result emphasizes that improving productivity not only reduces the required investment by the firm but also dramatically boosts the firm’s profit potential.
@@
The change rate of alpha starts at a high level (54.47%) when $\beta=0.40$ and drops to $-20.96\%$ at $\beta=1.00$, showing that better productivity substantially reduces the need for the firm to increase the level of investment. In terms of profitability, the profit change rate for the firm ranges widely, from $-79.50\%$ to a remarkable $131.44\%$. This further illustrates how vital firm’s productivity improvements are for enhancing its profitability. Essentially, better productivity allows firms to generate substantially higher profits with a reduced level of investment, making this an essential factor in optimizing their financial strategies of the firm.
Carbon Price Per Unit $(B)$ :

### v2

@@ A Total Asset
II Carbon Intensity
Table 1: Variables of the Model
@@ 2.3 Optimal Transition Investment Ratio without Government Intervention
The transition towards a low-carbon economy presents a complex challenge for enterprises, acting as a double-edged sword with respect to both arising opportunities and costs. While embracing low-carbon transformation holds the promise of environmental sustainability and enhanced competitiveness, it also entails a substantial amount of financial investments and operational adjustments for firms. For many enterprises, particularly those operating in carbon-intensive industries, the costs associated with transitioning to cleaner technologies and sustainable practices can be daunting. These costs may include investments in a renewable energy infrastructure, efficiency upgrades, and emission reduction measures, all of which require a substantial amount of capital expenditure and major operational restructuring in order to embark on the transition path.
As a result, some companies may hesitate to fully commit to a low-carbon transformation, fearing the resulting financial burden and potential disruptions to their existing business models. This reluctant proclivity to adopt green practices could undermine a firm’s progress towards sustainability goals and exacerbate environmental challenges. In such cases, the role of government intervention becomes relevant in incentivizing and regulating enterprises’ low-carbon transformation efforts.
@@
Hence, this section aims at analyzing and delineating the optimal low-carbon transition investment strategies for companies, considering both scenarios with and without external government intervention. By comparing these two approaches, we hope to elucidate the efficacy and implications of different pathways toward sustainable transformation within the corporate landscape and to provide valuable insights and suggestions for policymakers, such as government entities, to better support and incentivize effective low-carbon strategies in the corporate sector.
@@
Maximizing profit is a prime objective of a company. This strategic action emanates from the basic principle that profit arises from the difference between total revenue and total costs. Therefore, to effectively navigate and optimize profit margins, it becomes indispensable to dissect the constituents of both total revenue and total costs of the company. In our model, total revenue is represented by the unit selling price per unit of the product (denoted as $p$) multiplied by the total production volume. The total production volume is represented by a stylized sigmoid function, $2\cdot\text{Sigmoid}(\alpha k)\cdot[(1-\alpha)\cdot A]^{\beta}$, where $\operatorname{Sigmoid}(\alpha k)=\frac{1}{1+e^{-\alpha k}}$. This stylized formulation integrates several key variables: the transition investment ratio ($\alpha$), indicating the allocation of total assets towards low-carbon initiatives; the enterprise low-carbon production efficiency coefficient ($k$), denoting the effectiveness of transitioning to cleaner technologies; and the original productivity coefficient ($\beta$), reflecting the baseline production efficiency. The sigmoid function encapsulates the joint impact of transition investments and the firm’s production efficiency on production volume, ensuring a realistic representation of the firm’s saturation effects. Meanwhile, the term $(1-\alpha)\cdot A$ captures the portion of the firm’s total assets available for production activities, adjusted by the original productivity coefficient $\beta$ to reflect the company’s inherent efficiency before investing in the low-carbon transition. Thus, the term $(1-\alpha)\cdot A]^{\beta}$ represents the firm’s total production volume before investing in a low-carbon transition. Simultaneously, the firm allocates a portion $\alpha$ of its total assets $A$ towards the low-carbon transition investment. This investment is expected to enhance production efficiency, leading to an overall increase in production volume by a factor of $2\cdot\operatorname{Sigmoid}(\alpha k)$. By combining these various elements, the formula provides a greater understanding of how strategic decisions regarding transition investments and production efficiency influence the overall production capacity of the company.
@@
The design of this production function originates from a Cobb-Douglas production function, which is a widely used model in Economics to link a set of inputs (typically capital and labor) to an output in a production process. Its form with constant returns to scale is $Q=A\cdot K^{\beta}\cdot L^{1-\beta}$, where $Q$ is a total output, $A$ is a scaling or efficiency parameter, $K$ is a capital input, $L$ is a labor input, and $\beta$ is the elasticity of output with respect to capital. In our model, the term $(1-\alpha)\cdot A]^{\beta}$ corresponds to $K^{\beta}$. The sigmoid function, multiplied by 2, acts as an adaptive transformation of the efficiency parameter $A$ in the Cobb-Douglas production function, reflecting the transition investment’s impact on production efficiency. Unlike the traditional Cobb-Douglas function, which considers both capital and labor inputs, our model excludes the labor component due to the specific context of low-carbon investment decisions. This modification is intended to ensure that the production process incorporates diminishing marginal returns to transition investments ( $\alpha$ ) and captures the saturation effects as efficiency reaches its practical limits. Notably, we introduce a scaling factor of ‘2’ at the beginning of the sigmoid function. This adjustment is made primarily to ensure that the sigmoid’s baseline value remains at least 0.5 when the input ($\alpha k$) exceeds 0 in our model. By multiplying the sigmoid by ‘2’, we normalize its baseline rate to 1, resulting in a more pronounced increase above 1 as the values of $\alpha$ (Transition Investment Ratio) or $k$ (Enterprise Low-carbon Production Efficiency Coefficient) rise. This modification aligns with our prior expectations about this relationship.
@@
Our results are expected to hold broadly with alternative forms of the production functions exhibitting concavity and diminishing marginal returns, which are the main properties captured by the sigmoid-based function. Once such production function is a Constant Elasticity of Substitution (CES). The trade-offs between transition investments and production efficiency, as well as the insights on profitability, depend on these shared characteristics and are not restricted to the functional form that we use in our study. However, the CES function, which allows for variable elasticity of substitution between inputs, could confer an additional degree of flexibility and provide further insights into how substitutability between capital and efficiency affects production outcomes.
Given the above description of the model, we define ‘Revenue’ as:
Revenue = $=\text{ Selling Price per unit }\times\text{ Unit Amount }$ (2.1)
@@
Next, we address the ‘Cost’ component of the model, which comprises both the ‘Production Cost’ and the ‘Carbon Cost’. Denote ‘c’ as the production cost per unit. The production cost can then be expressed as ‘c’ multiplied by the unit amount (equivalent to the unit amount in the revenue part). Additionally, we incorporate the carbon cost in our model, which is calculated as the product of the ‘Carbon Intensity’, the unit amount, and the ‘Carbon Price’. Here, the ‘Carbon Intensity’ is defined as $\left(2-\frac{2}{1+e^{-\alpha k}}\right)$, which is a value set to range between 0 and 1. When $\alpha k$ is zero (indicating no transition investment), the carbon intensity reaches its highest level of 1. Conversely, as transition investment increases, the carbon intensity decreases, indicating a reduction in emissions due to greater transition efforts. As mentioned, we define ‘Carbon Intensity’ as the ‘Carbon Emission’ divided by the total production volume, aligning with the traditional understanding of ‘Carbon Intensity’ discussed earlier. Based on these considerations, we define the ‘Cost’ component in the model as:
$\text{Cost}=$ Production Cost + Carbon Cost (2.2)
@@
To determine the optimal value of $\alpha$, we take the first-order derivative of the total profit with respect to $\alpha$ and set the resulting equation to be eqiual to 0. This yields a necessary first-order condition (FOC) from the firm’s profit optimization problem:
@@
To determine whether the solution to the FOC yields a maximum value for the firm’s profit, we obtain a second-order derivative with respect to $\alpha$:
@@
We observe that the value is less than 0 when $\alpha$ and $\beta$ are both within the range of (0, 1). This leads us to conclude that the optimal value of $\alpha$ indeed maximizes the firm’s profit in keeping with our prior expectation.
As a first step, we try to understand how changes in specific variables affect the optimal transition investment ratio. For instance, we wish to determine how the optimal value of $\alpha$ adjusts when the carbon price ($B$) increases from 0 to 2. This analysis will elucidate the impact of carbon pricing on the optimal transition investment ratio for the firm. However, explicitly expressing $\alpha$ in terms of other variables analytically is highly intractable. The equation features exponential terms such as $e^{\alpha(-k)}$, which complicate the isolation of $\alpha$ since these terms do not simplify easily into a linear or polynomial form. Additionally, the equation includes a combination of linear and non-linear terms involving multiple variables ($A$, $B$, $c$, $p$, $\beta$, and $k$), further complicating the analysis. The presence of $\alpha$ in both the numerator and the denominator within the fractional expressions adds another layer of difficulty in the analysis. Therefore, below we will resort to numerical methods to assess how the optimal transition investment ratio of the firm varies in response to incremental parameter changes.
We conduct a numerical analysis of each critical variable in the model, which affects the optimal transition investment ratio ($\alpha$). Recall that the optimal transition investment ratio represents the proportion of firm’s financial resources allocated towards its low-carbon initiatives relative to its total investment. A key variable is a specific variable within the model that we systematically vary to observe its impact on the optimal $\alpha$. By focusing on one variable at a time, we can isolate its effect and gain deeper insights into its role within the model.
While holding all other variables constant in the model, we determine the optimal value of $\alpha$ for a range of values of a specific key variable. We then graphically represent the relationship between this key variable and the optimal $\alpha$. Additionally, we will present the maximum profit achievable for different values of the key variable. This approach allows us to visualize how changes in one variable influence both the optimal investment decision and the profitability of the company.
The baseline values of the variables in the model are provided in Table 2. These values are set within a reasonable range for illustrative purposes and can be adjusted by readers interested in exploring scenarios across different industries. Building on this foundation, we will systematically vary one variable at a time to different values and analyze the resultant changes in the value of $\alpha$.
@@
pcp-c Selling Profit per unit 2
@@ A Total Asset 100
Table 2: Baseline Values of Variables in the Model
From Figure 1 to Figure 4, the left plot illustrates how changes in the key variable in the model affect the optimal transition investment ratio, with the goal of maximizing profit under each scenario. The right plot shows the corresponding maximum profit achieved under the optimal transition investment as the key variable changes. We will explain each figure in detail below.
Relationship Between Profit Margin (pcp-c) and Optimal Alpha (Figure 1). Figure 1 reveals a clear pattern where the optimal value of $\alpha$ for the firm remains high and constant initially, even when the selling profit per unit is very low or negative, suggesting a strategic focus on substantial investments in machinery upgrades or efficiency improvements during periods of low profitability. This high investment level, indicated by a stable high value of $\alpha$, which is likely to reflect a counterintuitive but proactive approach: companies invest heavily during financially challenging times to enhance production efficiency and reduce operational costs, aiming at setting the groundwork for future competitiveness and sustainability.
As the selling profit per unit begins to increase for the firm, the optimal value of $\alpha$ sharply declines, indicating that further heavy investments in transformation by the firm are no longer necessary. The decrease in the value of $\alpha$ suggests that the earlier investments by the firm have begun to yield improved operational efficiencies and higher profits, reducing the need for a continued high expenditure in transformations. This strategic shift emphasizes a dynamic financial strategy where initial heavy investments by the firm during low-profit periods are scaled back as profitability improves, aligning the investment levels of the firm with the diminishing marginal benefits of further transformations.
Figure 1: Optimal Level of Alpha and Profit Graph w.r.t. pcp-c
Effect of Increasing Carbon Prices ($B$) on Optimal Alpha (Figure 2). As shown in Figure 2, when the carbon pricing variable $B$ increases, it raises operational costs for the firm due to higher carbon taxes. In response, the optimal investment ratio $\alpha$ also increases, as shown in the left plot. This adjustment indicates a strategic move by the firm to allocate more resources towards low-carbon transition initiatives, which enhances production efficiency and reduces carbon emissions, thereby helping to offset the financial burden imposed by the carbon tax.
@@
Impact of Efficiency Parameter ($k$) on Optimal Alpha and Profit (Figure 3). Combining insights from both plots in Figure 3, we can observe that as the efficiency parameter $k$ increases, profit consistently rises, while the optimal investment ratio $\alpha$ peaks at around $k\approx 3.1$ and then gradually declines. This suggests that although higher $k$ values are beneficial for the company’s profitability, it does not require continuously increasing the investment ratio in low-carbon initiatives. Initially, when $k$ is low, the marginal benefit from efficiency gains is high, encouraging a higher allocation to low-carbon investments. However, as $k$ grows, the marginal efficiency gains diminish, making it less necessary to maintain a high $\alpha$. At higher $k$ values, profits continue to increase due to the enhanced production efficiency achieved by prior investments, allowing the company to reduce the investment ratio $\alpha$ while still benefiting from increased profits.
@@
Influence of Original Productivity Coefficient ($\beta$) on Optimal Alpha and Profit (Figure 4). The plots in Figure 4 demonstrate that as the Original Productivity Coefficient ($\beta$) increases, profit grows exponentially while the optimal transition investment ratio ($\alpha$) declines. A higher $\beta$ indicates strong baseline productivity, reducing the need for extensive low-carbon investments since the firm’s production efficiency is already high. This makes the additional benefits from low-carbon investments less pronounced, allowing the firm to allocate fewer resources to $\alpha$. The exponential growth in profit with rising $\beta$ highlights the impact of inherent productivity on profitability, as firms with high baseline efficiency can achieve substantial profits without heavy investment in low-carbon initiatives. The decrease in optimal $\alpha$ reflects diminishing marginal returns on low-carbon investments for high-$\beta$ firms, where the benefits of further investment in $\alpha$ are minimal.
@@
The sensitivity analysis is conducted to evaluate the impact of variations in critical variables on the optimal $\alpha$, the maximum objective value, and the profit change rate. The variables scrutinized include the enterprise low-carbon production efficiency coefficient ($k$), production cost per unit ($c$), original productivity coefficient ($\beta$), and carbon price per unit ($B$). Adjustments for each parameter are performed at ±10% and ±50% based on the baseline values in Table 2, except for $\beta$, where the upper limit is constrained to 1.00 due to inherent limitations. The results are presented in the Table 3 and the analysis of each variable and parameter is reported below.
@@
Profit Change Rate (%) 60.249582 11.842530 0.000000 -11.69191 -56.05747
@@
Alpha Change Rate (%) 54.473780 9.273146 0.000000 -8.725114 -20.955217
@@
Profit Change Rate (%) 26.669326 4.834262 0.000000 -4.600030 -20.871997
@@
Enterprise Low-Carbon Production Efficiency Coefficient $(k)$ . The sensitivity analysis for $k$ reveals that as $k$ increases from 1.0 to 3.0, the optimal $\alpha$ rises from 0.14 to 0.41, which indicates that improving a firm’s low-carbon production efficiency encourages a larger investment by the firm in low-carbon technologies. This trend is further reflected in the maximum objective value, which grows from 40.39 to 62.52, showing that better firm’s efficiency leads to higher profitability for the firm. These results highlight a strong connection between firm’s low-carbon production efficiency and its profitability.
Moreover, the $\alpha$ change rate starts with a large negative value (-62.77%) when $k=1.0$ and makes a transition to a moderately positive change of $9.14\%$ by $k=3.0$. Similarly, the firm’s profit change rate shifts from a notable loss ($19.92\%$) at $k=1.0$ to a relatively large gain $(23.96\%$) at $k=3.0$. This indicates that as efficiency improves, firms not only increase their levels of investment in cleaner technologies but also experience greater profitability. These results underline the importance of improving low-carbon production efficiency, as it can serve as a key driver for both firm’s optimal investment and its profit maximization.
@@
Production Cost Per Unit (c). As $c$ increases from 0.80 to 2.40, the optimal $\alpha$ rises from 0.30 to 0.52, suggesting that firms must invest more to counterbalance higher production costs. However, this comes at the cost of its profitability, as the maximum objective value drops sharply from 80.82 to 22.16, showing that rising production costs significantly undermine a company’s profitability potential.
@@
The $\alpha$ change rate starts at $19.88\%$ for $c=0.80$, meaning that lower production costs lead to less investment in low-carbon technology. As $c$ rises, the $\alpha$ change rate becomes positive (38.13%), showing that companies need to increase their investments to cope with higher costs. Profit change rates, meanwhile, vary drastically, reaching a peak at $60.25\%$ for $c=0.80$, followed by a steep decline to $56.06\%$ at $c=2.40$. This result illustrates the sensitivity of firm’s profitability to changes in its production costs and highlights the point that controlling production costs is crucial for firms whose goal is to sustain profits while investing in low-carbon solutions. If its production costs rise unchecked, the firm may struggle to maintain its profitability, even with optimal investments.
@@
Original Productivity Coefficient $(\beta)$ . When it comes to the value of $\beta$, the sensitivity analysis overall shows a unique pattern: as $\beta$ increases from 0.40 to 1.00, the optimal $\alpha$ decreases from 0.58 to 0.30. This indicates that as its productivity improves, the firm requires less investment in low-carbon technologies to maintain profitability. Simultaneously, the maximum objective value surges from 10.34 to 116.73, highlighting the substantial impact of productivity on overall profitability of the firm. This result emphasizes the fact that improving productivity not only reduces the required investment by the firm but also dramatically boosts the firm’s profit potential.
@@
The change rate of $\alpha$ starts at a high level (54.47%) when $\beta=0.40$ and drops to $20.96\%$ at $\beta=1.00$, showing that better productivity substantially reduces the need for the firm to increase the level of investment. In terms of profitability, the profit change rate for the firm ranges widely, from $79.50\%$ to a remarkable $131.44\%$. This further illustrates how vital firm’s productivity improvements are for enhancing its profitability. Essentially, better productivity allows firms to generate substantially higher profits with a reduced level of investment, making this an essential factor in optimizing their financial strategies of the firm.